OpenStudy (anonymous):

combining functions x^2-1 ------ x^2+x

OpenStudy (anonymous):

\[\frac{ \left( x-1 \right) \left( x+1 \right)}{ x \left( x+1 \right) }=\frac{ \left( x-1 \right) }{ x }\] something like that

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

x cannot be -1 or zero

OpenStudy (anonymous):

for x = -1 the answer is zero 0/-1

OpenStudy (anonymous):

(-1)^2 + (-1) = 0

OpenStudy (anonymous):

yes 0/-1 is 0 >> that it can be x=-1 but x=0 >> f(x) does not exist

OpenStudy (anonymous):

For x is 0 or -1 the original expression has a denominator of 0 (check it)

OpenStudy (anonymous):

I mean for (x-1)/x ><"

OpenStudy (anonymous):

That's what I am telling you, you cannot just cancel the x+1....

OpenStudy (anonymous):

OK :)

OpenStudy (anonymous):

I do not know the objective of this prob. to find ???

OpenStudy (anonymous):

Or, you can, only if you specify that x is not 0 or -1

OpenStudy (anonymous):

I don't know either, what "combining" means?

OpenStudy (anonymous):

It's a division problem

OpenStudy (anonymous):

Pinguine has given the answer but you should also specify that the original expression is undefined for x = 0 or -1